Solve 2x + 3y = 11 and 2x – 4y = – 24 and hence find the value of ‘m’ for which y = mx + 3. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2x + 3y = 11…………………………..(I)
2x – 4y = -24………………………… (II)
From equation (II)
we get
x = (11-3y)/2 ………………….(III)
Substituting the value of x in equation (II),
we get
2(11-3y)/2 – 4y = 24
11 – 3y – 4y = -24
-7y = -35
y = 5……………………………………..(IV)

Putting the value of y in equation (III),
we get
x = (11-3×5)/2 = -4/2 = -2
Hence
x = -2,
y = 5
Also
y = mx + 3
5 = -2m +3
-2m = 2
m = -1
Therefore the value of m is -1.



Thanks